OpenStudy (anonymous):

what's the square root of 6x^3multiplied by the square root of 18x^4

OpenStudy (espex):

\[\sqrt{6x^{3}}\times\sqrt{18x^{4}}\rightarrow\sqrt{6\times18 \times x^{3}\times x^{4}}\rightarrow\sqrt{6\times6\times3\times x^{3}\times x^{2}\times x^{2}}\]